The dimensions of energy and torque respectively are
A. ML^2T^-2 and ML^2T^-2
B. MLT^2 and ML^2T^-2
C. ML^2T^-2 and MLT^-2
D. MLT^-2 and MLT^-2

Grade:Upto college level

1 Answers

Apoorva Arora IIT Roorkee
9 years ago
Energy = Force * distance=Mass*acc*distance=M*LT-2*L=ML2T-2
Torque =Force * distance=Mass*acc*distance=M*LT-2*L=ML2T-2
Thanks and Regards
Apoorva Arora
IIT Roorkee
askIITians Faculty